OpenStudy (anonymous):

a tension force of 2.50 N acts horizontally on a 2.00 kg block. The block accelerates at 0.750 m/s2. What is the force of kinetic friction?

OpenStudy (anonymous):

F = m a 0.750 m/s2 = a = F/m = (2.50 N - friction) / (2 kg) The rest is yours!